SHerry has 23 coins in his ashtray. There are nickels, dimes and quarters. If he has half as many quarters as nickels and a total of $2.65, how many of each type of coin does he have?

Answer by ikleyn(52786)

.

n  quarters;  2n nickels and the rest  23-n-2n = 23-3n  are dimes.


The total money equation is then


    25n + 5*(2n) + 10*(23-3n) = 265  cents,   or


    25n + 10n + 230 - 30n = 265

    5n = 265 - 230 = 35

     n = 35/5 = 7.


ANSWER.  7 quarters;  14 nickels and the rest  are  23-7-14 = 2 dimes.

Solved.

---------------

The miracle is that the problem is solved using only one unknown and one single equation.